b)Mr.Ali has an income of Rs.30,000 and he purchases 100 units of X-GOOD. His income decreases by 20% and now he can purchase 90 units of X-GOOD.Calculate Calculate elasticity and also mention the type of elasticity.
1
Expert's answer
2020-05-03T16:58:24-0400

The income elasticity of demand is computed as:



"E_Y = \dfrac{\%\Delta Q}{\%\Delta Y}"

The percentage change in demand is:



"\%\Delta Q = \dfrac{90 - 100}{100} \times 100 = -10\%"

If the income decreases by 20%, the income elasticity of demand is equal to:



"E_Y = \dfrac{-10\%}{-20\%}"

"\boxed{\color{red}{E_Y = 0.5}}"